OpenStudy (anonymous):

Complete the equations of the system in slope-intercept form. Use a decimal for the slope if necessary. Line 1 x y 0 3 1 5 Line 2 x y –1 1 –2 –1 Line 1: y = x + Line 2: y = x +

OpenStudy (anonymous):

@MAli13chineta

OpenStudy (anonymous):

im not sure sorry @PlanSlam @wio @TheSmartOne

OpenStudy (anonymous):

Okay, so the thing to understand about lines, is that the change in \(y\) and change in \(x\) are always proportional.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

For line 1, the change in \(y\) is given by \(y_2-y_1=5-3=2\). The change in \(x\) is given by \(x_2-x_1=1-0=1\). You get the slope by dividing change in \(y\) by change in \(x\).

OpenStudy (anonymous):

For line 1, the slope is \(m=2/1=2\)

OpenStudy (anonymous):

so y=2 x+2?

OpenStudy (anonymous):

The \(y\) intercept is what you get when you let \(x=0\).

OpenStudy (anonymous):

Line 1 already tells us that when \(x=0\), we have \(y=3\)

OpenStudy (anonymous):

So with a slope of \(2\) and \(y\) intercept of \(3\), our line is:\[ y=2x+3 \]

OpenStudy (anonymous):

In general, if you have two points \((x_1,y_1)\) and \((x_2,y_2)\), you can find the equation for the line by solving for \(y\): \[ \frac{y-y_1}{x-x_1}=\frac{y_2-y_1}{x_2-x_1} \]So for line 2, we would start with:\[ \frac{y-(1)}{x-(-1)}=\frac{(-1)-(1)}{(-2)-(-1)} \]

OpenStudy (anonymous):

Using algebra on:\[ \frac{y-y_1}{x-x_1}=\frac{y_2-y_1}{x_2-x_1} \]to solve for \(y\) would give:\[ y = \frac{y_2-y_1}{x_2-x_1}(x-x_1)+y_1 \]

OpenStudy (anonymous):

Sorry @wio my connection went fuzzy and had to refresh